Use the data entry method in your scientific calculator to enter the following measurements.
x 1 2 3 4 5 6
y 5.5 4.4 4.3 3.6 3.0 2.5
Recall the proper memories to find the correlation coefficient, r. (Round your answer to four decimal places.)
r =
Recall the proper memories to find the y-intercept, a, of the line. (Round your answer to three decimal places.)
a =
Recall the proper memories to find the slope, b, of the line. (Round your answer to three decimal places.)
b =

Answer by Theo(13342)

Recall the proper memories to find the correlation coefficient, r. (Round your answer to four decimal places.)

results of the linear regression online calculator at https://stats.blue/Stats_Suite/correlation_regression_calculator.html are shown below.



answers to your questions are below:

the correlation coefficient = -.985217
the slope = -.568571
the y-intercept = 5.873333

note that this calculator has the equation in y = ax + b format.
a is the slope
b is the y-intercept.

note that other calculators can have the equation in y = a + bx format.
a is the y-intercept
b is the slope

be careful what format the calculator that you use is in.

my calculator (ti-84 plus) gives you the option as to which format you want to display.

the online calculator that i used shows is in the y = ax + b format.

the correlation coefficient is always shown as a coefficient to the variable.
the y-intercept is always shown separately as a constant.